### 26 Apr 2026 v2

# When Forecast Accuracy Fails: Rank Correlation and Decision Quality in Multi-Market Battery Storage Optimization

Abstract
[…]
Keywords: battery energy storage; multi-market optimization; forecast evaluation; Kendall tau; rolling intrinsic; frequency containment reserve; Swiss balancing market
1 Introduction
Grid-scale battery energy storage systems (BESS) have become central to European electricity market operations, with prequalified FCR capacity in Germany growing by 180 MW in 2024 alone, and aFRR prequalifications increasing by over 400% year-on-year (Regelleistung.net, 2024). As these assets compete across multiple simultaneous markets — frequency containment reserve (FCR), automatic frequency restoration reserve (aFRR), day-ahead (DA), and continuous intraday (XBID) — their commercial viability increasingly depends on automated, forecast-driven optimization systems. However, the effectiveness of such systems depends critically on how forecast information is translated into economic decisions.
Recent work has advanced multi-market BESS optimization substantially. Schaurecker et al. (2025) demonstrate that a dynamic programming formulation of the rolling intrinsic strategy captures near-optimal intraday trading profits. Seifert et al. (2024) develop a stochastic multi-market bidding framework based on SDDP for coordinated participation across FCR, day-ahead, and intraday markets, finding that coordination does not necessarily improve revenues and that FCR often dominates optimal strategies. Hornek and others (2025) compare forecast-driven and perfect-foresight strategies on the continuous intraday market, finding that a forecast-driven system captures approximately 90% of perfect-foresight revenue (achieving revenues only about 11% below perfect foresight).
A parallel literature has begun to examine the mismatch between forecast accuracy and economic value. Uniejewski and others (2025) demonstrate that MAE and RMSE are only weakly correlated with BESS profits in the context of day-ahead electricity markets and BESS arbitrage, capturing only a limited dimension of forecast performance. Vanderschueren et al. (2022) show more broadly that the decision strategy dominates the training strategy in cost-sensitive optimization, and Smets et al. (2024) propose decision-focused training for BESS market participation. Building on the foundational work of Morales et al. (2013) on forecast value in energy systems, this literature increasingly recognizes that statistical accuracy and economic value diverge. Several of these contributions are recent preprints that have not yet completed peer review; we include them because they represent the current frontier of this rapidly developing literature.
Yet two gaps remain. First, the continuous intraday market is structurally more ordinal than day-ahead auctions: the DP dispatch problem requires only the relative ranking of future prices, not their absolute levels. No existing study has exploited this structural property to identify a formal metric that predicts decision quality in intraday BESS trading. Second, the literature is almost exclusively focused on the German market; the Swiss balancing market—structurally distinct due to its hydro-dominated reserve provision—remains entirely unexplored. As a result, it remains unclear whether improvements in statistical forecast accuracy translate into improved economic outcomes in real trading systems, and how market-specific structural drivers shape revenue in non-German markets.
This paper makes three contributions. First, and most importantly, we identify a formal tau-sufficiency region ($\tau^{*}\approx 0.85$$0.95$) in the continuous intraday market, beyond which value capture saturates near its theoretical maximum regardless of further improvements in forecast accuracy. While Uniejewski and others (2025) demonstrate that rank-based metrics better predict BESS profits than MAE in day-ahead markets, we identify a quantified threshold—not merely a correlation—and show that this threshold is stable across market regimes, volatility levels, and subperiods. The structural explanation is that the DP dispatch problem is inherently ordinal: it depends on the relative ordering of prices rather than their absolute values (Morales et al., 2013).
Second, we demonstrate that revenue in multi-market BESS systems is primarily driven by capacity allocation rather than intraday forecast quality. Under reserve market constraints (PICASSO compliance), the residual SoC budget allocated to XBID yields revenues substantially lower than FCR per MW committed—not because intraday arbitrage is structurally less profitable, but because co-participation in reserve markets physically limits the available intraday cycling capacity. We find no evidence of systematic misallocation of capacity across markets within the scope of our model and dataset under realistic operating conditions.
Third, we provide the first empirical analysis of the Swiss balancing market using real Swissgrid data, showing that hydrological conditions are strongly associated with balancing prices and battery revenues, highlighting the importance of market-specific structural drivers absent from existing German-focused literature.
The remainder of this paper is organized as follows. Section 2 describes the market architecture and data. Section 3 presents the hierarchical optimization system and the forecast evaluation framework, including formal definitions of the Value Capture Ratio and tau-sufficiency. Section 4 reports empirical results in order of contribution strength. Section 5 discusses theoretical and practical implications, and Section 6 concludes.
2 Market Structure and Data
2.1 Market Architecture
Battery storage systems in Central Europe participate in a cascade of sequential markets, each with distinct gate closures, clearing mechanisms, and revenue structures. We consider four markets in Germany and their Swiss equivalents.
The FCR is procured through daily auctions covering six four-hour blocks per day, operated jointly by twelve European TSOs under the FCR Cooperation framework. Participants submit capacity bids; accepted bids receive a pay-as-bid capacity payment with no separate energy settlement. In Germany, the minimum bid size is 1 MW and bids must be symmetric (equal upward and downward capacity). Gate closure is D-1 at 08:00 CET.
The aFRR is procured through daily capacity auctions at D-1 09:00 CET, separately for upward (aFRR++) and downward (aFRR-) products, under the ALPACA cooperation framework for Germany and Austria. Accepted capacity bids receive a capacity payment; energy activations by the TSO are settled at the PICASSO platform price. Under PICASSO compliance rules, providers must maintain a SoC buffer sufficient to sustain full activation for a minimum of 30 minutes (FCR) and 60 minutes (aFRR), limiting the effective capacity available for other markets.
The Day-Ahead (DA) market clears at D-1 12:00 CET through the EPEX SPOT single price auction under the Single Day-Ahead Coupling (SDAC). Hourly products are cleared at a uniform price across the coupled zone.
The continuous Intraday (XBID) market opens at D-1 15:00 CET and closes 30 minutes before physical delivery in Germany. Trading occurs continuously via a shared European Limit Order Book until gate closure.
The Swiss market parallels Germany with two structural differences. First, Swissgrid procures Secondary Reserve (SRL++/SRL-) through weekly tenders rather than daily auctions, fixing prices for the entire week. Second, the XBID gate closure in Switzerland is 60 minutes before delivery, reducing the effective intraday trading window. The Swiss reserve pool is approximately 400 MW, supplied predominantly by flexible hydroelectric generation. When reservoir levels are anomalously high relative to seasonal norms—as during the 2022–2023 Speicherkrise—TSOs require additional downward flexibility to absorb excess hydraulic generation, driving SRL- prices to extreme levels. This mechanism is the subject of our natural experiment in Section 4.3.
These market characteristics determine both the timing and the informational requirements of forecasting and decision-making across layers.
2.2 Dataset
Table 1 summarizes the data sources used in this study. All data are publicly available and aligned to a common 15-minute timeline; processing is designed to avoid any form of look-ahead bias. The dataset covers January 2020 through December 2025, yielding six full calendar years across two distinct market regimes: the energy crisis of 2021–2022 and the post-crisis normalization of 2023–2025.
Table 1: Summary of data sources used in this study. All data are publicly available.
Source Variable Market Period Freq.
Regelleistung.net FCR/aFRR capacity prices DE 2020–2025 4h
SMARD Day-ahead prices DE 2020–2025 15min
SMARD XBID prices DE 2020–2025 15min
SMARD Wind, solar, load DE 2020–2025 15min
Swissgrid SRL++/SRL- tender prices CH 2015–2025 weekly
Swissgrid aFRR activation & prices CH 2020–2025 15min
Swissgrid Hydro reservoir levels CH 2020–2025 weekly
ENTSO-E Day-ahead prices CH, AT, FR 2020–2025 hourly
ENTSO-E Generation by source DE 2020–2025 15min
Open-Meteo ERA5 Temperature, wind DE/CH 2020–2025 hourly
ICE / yfinance Gas TTF, CO2 EUA EU 2020–2025 daily
German FCR and aFRR capacity prices are obtained from Regelleistung.net at four-hour resolution. Day-ahead and intraday prices for Germany are obtained from the SMARD platform operated by the Federal Network Agency (Bundesnetzagentur) at 15-minute resolution. Swiss SRL tender prices are obtained from Swissgrid’s public data portal at weekly resolution; Swiss aFRR activation data and reservoir levels from the same source at 15-minute and weekly resolution respectively. Day-ahead prices for Switzerland, Austria, and France are from the ENTSO-E Transparency Platform. Meteorological variables (temperature, wind speed, solar irradiance) are from the Open-Meteo ERA5 reanalysis at hourly resolution. Gas (TTF) and carbon (EUA) prices are from ICE and Yahoo Finance at daily resolution.
2.3 Data Quality and Methodological Notes
Several corrections are required to ensure economic consistency and avoid bias in the simulation framework.
Volume-weighted average correction for Swiss SRL prices.
Raw Swissgrid tender data reports arithmetic mean prices across accepted bids. We replace these with volume-weighted average prices (VWA), which better reflect the effective clearing price faced by a price-taking participant. This correction reduces the estimated SRL- price by 13–18% in normal market conditions and has negligible effect during crisis periods when the spread between accepted bids is small.
Leakage removal in LGBM forecasting.
Early versions of the feature set included contemporaneous FCR clearing prices as a predictor for DA price forecasting. This constitutes look-ahead bias, as FCR clears at 08:00 while DA clears at 12:00 on the same day. Removing FCR prices from the training set increases the honest DA MAPE from 11% (inflated) to 39% (realistic), consistent with benchmarks reported for the German market in the literature.
Pay-as-bid bid mechanics.
FCR and aFRR capacity markets in Germany use pay-as-bid pricing: each accepted bidder receives their own submitted price. Bid prices in our model are set to the $Q_{40}$ quantile of the rolling historical price distribution (static strategy) or adapted by the HMM++SAC regime detection layer (adaptive strategy). Acceptance probability is modeled as $\Pr[\text{accepted}]=\Pr[\text{bid}\leq p_{\text{clear}}]$ using the empirical clearing price distribution from Regelleistung.net. This formulation ensures that expected revenue reflects both bid level and acceptance probability. Together, these data and corrections provide the empirical foundation for the hierarchical optimization system described in Section 3.
3 Methodology
3.1 System Architecture
The optimization system follows a hierarchical three-layer architecture, with each layer operating at a different temporal resolution and solving a distinct subproblem. The layered structure is modular: each component can be replaced without affecting the overall framework. A single physical SoC trajectory is shared across all layers, ensuring that decisions at each level remain physically consistent with the battery’s actual energy state.
Layer 1 — Weekly capacity allocation.
Layer 1 determines how the battery’s limited power capacity is allocated across markets with different revenue structures. At the beginning of each week, a stochastic linear program allocates capacity across FCR, aFRR, and DA for the following seven days, maximizing expected revenue over 30 price scenarios drawn from the historical price distribution, subject to PICASSO compliance constraints. The PICASSO buffer reserves a fraction of the battery’s energy capacity to cover worst-case FCR and aFRR activations:
$\text{SoC}_{\min}=\frac{P_{\text{FCR}}\cdot 0.5\,\text{h}+P_{\text{aFRR}}\cdot 1.0\,\text{h}}{S_{\max}}$ (1)
where $P_{\text{FCR}}$ and $P_{\text{aFRR}}$ are the allocated MW in each reserve market and $S_{\max}=10\,\text{MWh}$. The output of Layer 1 is a weekly MW allocation schedule and a bid price for each reserve product, generated by the HMM++SAC adaptive bidding strategy described below.
Layer 2 — Daily schedule optimization.
Layer 2 plays a supporting role by smoothing inter-day SoC dynamics. After the Day-Ahead market clears at D-1 12:00 CET, a rolling Model Predictive Control LP optimizes the hourly charge/discharge schedule over a five-day horizon, given the realized DA prices and MW allocations from Layer 1. LGBM-based price forecasts for D++1 through D++4 are used with exponentially declining weights to reflect forecast uncertainty.
Layer 3 — Intraday dispatch.
Every 15 minutes, a Dynamic Programming Rolling Intrinsic algorithm re-optimizes the XBID trading schedule over the remaining delivery slots of the current day, using the updated XBID price forecast as input. The DP solves:
$V_{t}=\max_{\delta\in[-F_{\max},\,F_{\max}]}\Bigl[\hat{p}_{t}\cdot\delta\cdot\Delta t+V_{t+1}\!\left(\text{SoC}_{t}+\delta\cdot\eta\cdot\Delta t\right)\Bigr]$ (2)
where $\hat{p}_{t}$ is the forecast XBID price for slot $t$, $\delta$ is the charge/discharge decision (positive == discharge), $\Delta t=15\,\text{min}$, and $\eta$ is the round-trip efficiency. The SoC is constrained to $\text{SoC}_{\min},\,\text{SoC}_{\max}]$ inherited from Layers 1 and 2.
The key structural property of this formulation is that the DP decision depends primarily on the relative ordering of future prices $\hat{p}_{t}$, not their absolute levels: a price sequence and any monotone transformation thereof yield identical charge/discharge decisions. This property implies that absolute price errors do not affect optimal decisions as long as the ranking is preserved—a result we formalize and quantify in Section 4.1.
Adaptive bidding — HMM and SAC.
The static $Q_{40}$ bid strategy performs poorly during regime transitions, because the bid price distribution shifts faster than the rolling window can adapt. Layer 1 incorporates a Hidden Markov Model (HMM) with four states (normal, crisis, post-crisis, low-volatility) trained on weekly FCR price dynamics. At each week, the HMM identifies the current market regime and a Soft Actor-Critic (SAC) agent selects the optimal bid percentile for that regime. The HMM++SAC system improved the FCR acceptance rate from 34% to 58% in the 2023 out-of-sample period, reflecting successful adaptation to the post-crisis normalization of FCR prices.
Battery parameters.
Power capacity $F_{\max}=10\,\text{MW}$, energy capacity $S_{\max}=10\,\text{MWh}$ (1-hour duration), round-trip efficiency $\eta_{\text{RT}}=0.9025$ ($\eta_{C}=\eta_{D}=0.95$), degradation cost $\nu_{\text{deg}}=4\,\text{EUR/MWh}$ applied via rainflow cycle counting with a depth-of-discharge exponent of 1.5. The system operates over 2020–2025 with walk-forward out-of-sample evaluation: model parameters are estimated on rolling training windows and applied to unseen data.
Modeling simplification — aFRR activations.
Energy activations under aFRR are modeled via expected value and SoC buffer constraints rather than fully dynamic real-time simulation. The PICASSO buffer in Equation (1) is designed to cover worst-case activation scenarios; actual intra-period SoC dynamics from activations are not simulated at 15-minute resolution. This assumption is conservative for Layer 3 revenue estimation: any dynamic activation that exceeds the buffer would reduce the available XBID capacity, lowering—not raising—our revenue estimates. The implications are discussed further in Section 5.4.
3.2 Forecast Evaluation Framework
The standard approach to evaluating electricity price forecasts relies on pointwise error metrics such as MAE and RMSE. While the value of forecasts can be theoretically formalized within stochastic optimization frameworks (Morales et al., 2013), recent empirical evidence suggests that improvements in standard forecast accuracy metrics do not necessarily translate into better operational decisions (Uniejewski and others, 2025). These metrics measure the average distance between predicted and realized prices across individual time slots, but are not aligned with the objective of the downstream optimization problem.
The limitation is not empirical but structural. A forecast can achieve low MAE by predicting prices close to their realized values in absolute terms, yet fail to reproduce the relative ordering of high-price and low-price slots within a day. For a battery operator, the economically relevant question is not “how close is the forecast to the realized price?” but rather “does the forecast correctly identify when to charge and when to discharge?” These are distinct questions with potentially opposite answers, as we demonstrate empirically in Section 4.1.
We formalize this distinction through three definitions.
Definition 1 (Decision Value).
Let $\pi^{*}$ denote the DP Rolling Intrinsic policy and $f$ a price forecast sequence. The decision value of $f$ is $V(\pi^{*},f)$, the revenue achieved by policy $\pi^{*}$ when using forecast $f$ as input. The decision value depends on the forecast only through the policy it induces: two forecasts that induce the same charge/discharge schedule yield identical decision values.
Definition 2 (Value Capture Ratio).
Given the oracle forecast $f_{\text{oracle}}$ (realized prices), the Value Capture Ratio is:
$\text{VCR}(f)=\frac{V(\pi^{*},f)}{V(\pi^{*},f_{\text{oracle}})}\in[0,1]$ (3)
The VCR normalizes revenue against the oracle benchmark, making it comparable across days and market conditions. A VCR of 1.0 means the forecast-driven strategy achieves the same revenue as perfect foresight; a VCR of 0.0 means no value is extracted.
Definition 3 (Ranking Inconsistency).
For a set of forecast models $\{f_{1},\ldots,f_{n}\}$, the pairwise ranking inconsistency between MAE and VCR is:
$\mathrm{RI}=\frac{|\{(i,j):\text{MAE}(f_{i})<\text{MAE}(f_{j})\text{ but }\text{VCR}(f_{i})<\text{VCR}(f_{j})\}|}{\binom{n}{2}}$ (4)
A high RI indicates that standard accuracy metrics are misleading guides for forecast model selection.
To measure the rank-preserving quality of a forecast independently of its level accuracy, we use the Kendall rank correlation coefficient:
$\tau_{K}(f)=\frac{(\text{concordant pairs})-(\text{discordant pairs})}{\binom{T}{2}}$ (5)
where a pair of slots $(i,j)$ is concordant if $\hat{p}_{i}>\hat{p}_{j}\Leftrightarrow p_{i}>p_{j}$ and discordant otherwise, and $T$ is the number of slots in the evaluation window. Unlike MAE, $\tau_{K}$ directly measures the preservation of pairwise price ordering, which is the sufficient statistic for the DP decision.
The connection to the DP formulation of Layer 3 is direct. As established in Section 3.1, the DP in Equation (2) depends primarily on the relative ordering of future prices, not their absolute levels. Therefore, any forecast that sufficiently preserves the ranking of prices induces a near-optimal policy. We define a forecast as tau-sufficient if $\tau_{K}(f)\geq\tau^{*}$ for a threshold $\tau^{*}$, and empirically identify $\tau^{*}\approx 0.85$$0.95$ in Section 4.1.
Our result is orthogonal to decision-focused learning (Smets et al., 2024), which retrains the forecaster to maximize downstream profit. We hold the decision policy fixed and ask what forecast quality is sufficient for near-optimal decisions. The tau-sufficiency threshold is a property of the decision problem structure—not of any particular forecasting method—and is therefore applicable regardless of how the forecast is generated.
3.3 Benchmark Suite
The benchmark suite is designed to span increasing levels of informational content, from no predictive structure to perfect foresight, allowing us to isolate whether improvements in statistical forecast accuracy translate into improvements in decision value.
Oracle uses realized XBID prices as forecast input, representing the upper bound on achievable revenue for a given SoC budget. By definition, $\text{VCR}(\text{oracle})=1.0$.
Persistence sets $\hat{p}_{t}=p_{t-96}$, replicating yesterday’s price profile. It can achieve relatively low MAE in stable market conditions as a naive baseline, but produces near-zero $\tau_{K}$ because it fails to capture intraday price dynamics. It serves as the critical test case for the MAEVCR divergence.
DA anchor uses the Day-Ahead clearing price for each hourly slot, replicated across 15-minute sub-slots, as the XBID forecast. Since DA prices are known at D-1 12:00 CET, this forecast requires no statistical model and is always available.
ML 1-step uses a LightGBM autoregressive model generating one-step-ahead forecasts via sequential rollout, trained on a rolling window of historical XBID prices, weather variables, DA prices, and calendar features. It achieves the highest $\tau_{K}$ among non-oracle forecasts.
Hybrid 8h blends ML 1-step forecasts within an 8-hour horizon with the DA anchor beyond, reflecting the principle that precise forecasts are only valuable within the time horizon where trades can be executed.
All forecasts are evaluated out-of-sample over 2023–2025 using the same DP Rolling Intrinsic policy, ensuring that differences in VCR are attributable solely to forecast quality. This design allows us to isolate whether improvements in statistical forecast accuracy translate into improvements in decision value.
3.4 Attribution Framework
To quantify the marginal contribution of each decision layer to total revenue, we construct an ablation study in which layers are added incrementally. A key methodological constraint governs the design: each configuration adds exactly one decision component, with all other parameters held fixed. This design isolates the marginal contribution of each decision layer to total revenue.
The configurations, ordered from minimal to full multi-market system, are:
  1. 1.
    Naive DA spread: charge during the six lowest DA price hours, discharge during the six highest. No reserve participation.
  2. 2.
    DP only: Layer 3 DP with full battery capacity available (no reserve obligations), using DA prices as a proxy for XBID in the absence of perfect foresight. VCR is computed relative to a standalone XBID oracle.
  3. 3.
    Layer 1 static ($Q_{40}$): FCR/aFRR allocation with static $Q_{40}$ bid, plus Layer 2 DA schedule and Layer 3 DP.
  4. 4.
    Layer 1 HMM: as above, replacing the static bid with the HMM-adaptive bid.
  5. 5.
    Full multi-market system: Layer 1 HMM++SAC, Layers 2 and 3.
  6. 6.
    Oracle (full multi-market system): perfect foresight on all markets simultaneously. Upper bound for the full system.
Each configuration is evaluated on the same 2023–2025 out-of-sample period. VCR is computed relative to a configuration-specific oracle to ensure that each configuration is evaluated against its own feasible upper bound, avoiding distortions due to differences in market participation. The standalone DP is compared to the standalone XBID oracle; all multi-market configurations are compared to the full-system oracle. The following section reports results in order of contribution strength, beginning with the forecast evaluation framework (Section 4.1), followed by the capacity allocation analysis (Section 4.2), and the Swiss natural experiment (Section 4.3).
3.5 Note on Synthetic Forecast Generation
The tau-sufficiency experiments reported in Section 4.1 use synthetic forecasts of controlled Kendall rank correlation, constructed via a linear interpolation between the realized price vector $y_{\text{true}}$ and an independent Gaussian noise vector $\varepsilon$:
$y_{\text{synth}}=\alpha\cdot y_{\text{true}}+(1-\alpha)\cdot\varepsilon$ (6)
with $\alpha$ calibrated by binary search to match each target $\tau_{K}$ value within tolerance $\pm 0.03$. This procedure achieves the target rank correlation by design but simultaneously attenuates the variance of the forecast proportionally to $(1-\alpha)$.
A robustness check using two alternative generation methods that preserve the marginal distribution of $y_{\text{true}}$ exactly—rank-perturbation, which permutes the indices of $y_{\text{true}}$ until the target $\tau_{K}$ is reached, and a Gaussian copula method with matching Spearman correlation—yields $\tau^{*}$ estimates approximately 0.100.100.120.12 higher than the alpha-interpolation method used in the main experiments. The mechanism is intuitive: a DP policy derived from variance-attenuated forecasts trades more conservatively, and is therefore less sensitive to ordering errors at the scale of the realized prices. Methods that preserve the full marginal variance penalize ordering errors at full price magnitude.
The qualitative findings of this paper—existence of a sufficiency threshold, rank correlation predicting VCR independently of pointwise error, stability across volatility regimes and market conditions—are robust across all three generation methods. The absolute threshold values $\tau^{*}\approx 0.85$$0.95$ reported herein should therefore be interpreted as a conservative lower bound under the variance-attenuating alpha-interpolation method; a measurement preserving the marginal distribution of prices yields $\tau^{*}\approx 0.95$$0.99$ for the same configuration. The structural claim that rank correlation dominates pointwise error in predicting decision value is unaffected by this calibration refinement.
The three generation methods considered—alpha-interpolation, rank-perturbation, and Gaussian copula—correspond to qualitatively different classes of synthetic forecast. Alpha-interpolation attenuates the marginal variance of $y_{\text{synth}}$ relative to $y_{\text{true}}$, a property shared by forecasts trained to minimize pointwise error losses such as MSE or MAE: the optimal point predictor (the conditional expectation $E[y\mid x]$ under MSE, the conditional median $\text{Med}(y\mid x)$ under MAE) has lower variance than $y$ whenever the conditioning information $x$ is informative. In this sense, alpha-interpolation captures a statistical signature that real ML forecasters typically exhibit regardless of the specific loss function employed, and the $\tau^{*}\approx 0.85$$0.95$ threshold reported here may be interpreted as the sufficiency level for this realistic class of attenuated forecasts. Rank-perturbation and Gaussian copula preserve the marginal distribution (exactly and asymptotically, respectively), and measure a purer ordinal sufficiency that is less tied to the statistical signature of pointwise-trained predictors. The slightly higher $\tau^{*}$ under these methods reflects this distinction. A fuller treatment of how $\tau^{*}$ depends on the class of forecasts considered lies outside the scope of this paper.
4 Results
4.1 The Tau-Sufficiency Region
We begin with the main empirical result of the paper. Figure 1 shows the relationship between Kendall $\tau_{K}$ and Value Capture Ratio, estimated from 24,000 DP simulations using synthetic forecasts of controlled rank correlation.
The relationship is highly non-linear. VCR grows steeply as $\tau_{K}$ increases from 0 to 0.70, with large marginal gains ($\approx$22 percentage points per 0.15 increase in $\tau_{K}$). Beyond $\tau_{K}\approx 0.85$, the curve enters a plateau: marginal VCR gains fall below 1.4 percentage points per step, and VCR stabilizes near 100% of oracle revenue. We define this plateau as the tau-sufficiency region.
Empirical Finding 1 (Tau-Sufficiency Region).
There exists an empirical threshold region $\tau^{*}\in[0.85,0.95]$ in Kendall rank correlation such that forecasts with $\tau_{K}(f)>\tau^{*}$ achieve $\text{VCR}(f)\geq 0.97$, even when absolute forecast errors remain non-negligible. Beyond this region, marginal improvements in forecast accuracy yield negligible economic value.
Figure 1: Empirical relationship between Kendall $\tau$ and Value Capture Ratio (VCR), estimated from 24,000 DP simulations with synthetic forecasts of controlled rank correlation. The tau-sufficiency region $[0.85,0.95]$ (shaded) marks the onset of VCR saturation near 100%. Points indicate the $\tau_{K}$ and VCR values of the five benchmark forecasts.
Table 2 reports the five benchmark forecasts evaluated on the 2023–2025 out-of-sample period. The results confirm the synthetic threshold finding. Persistence captures only 32.8% of oracle value, despite remaining a widely used baseline in electricity forecasting. By contrast, the ML 1-step forecast captures 99.2% of oracle value, and the Hybrid 8h forecast 99.7%. The key point is not a complete inversion of the benchmark ranking, but a strong misalignment between pointwise forecast error and economic usefulness: persistence may appear acceptable under simple statistical baselines, yet it destroys most of the value of the intraday dispatch problem, whereas forecasts with modest residual error already achieve near-oracle decisions once they enter the tau-sufficiency region.
Table 2: Forecast accuracy vs. decision quality for five benchmark forecasts. Kendall $\tau$ predicts VCR; MAE does not. Persistence achieves the highest MAE (30.1 EUR/MWh) and the lowest VCR (32.8%), confirming that statistical accuracy and decision quality move together in extreme cases. The more subtle finding is that beyond $\tau^{*}\approx 0.85$, further MAE reductions yield negligible VCR improvements: the Hybrid 8h forecast improves MAE by 53% over ML 1-step but gains only 0.5 percentage points in VCR. Formally, RI  = 0 across all five benchmarks, confirming that the misalignment is quantitative rather than ordinal. All results are out-of-sample (2023–2025, DE).
Forecast MAE $\tau_{K}$ VCR (%) MAE rk VCR rk
Oracle 0.00 1.000 100.0 1 1
Hybrid 8h 1.30 0.948 99.7 2 2
DA + Spread 2.22 0.928 99.5 3 3
ML 1-step 2.75 0.918 99.2 4 4
Persistence 30.1 0.000 32.8 5 5
MAE in EUR/MWh; $\tau_{K}$ = Kendall rank correlation with realized prices. Note: the key misalignment is not a ranking inversion across all benchmarks, but the large gap between Persistence ($\tau\approx 0$, VCR = 32.8%) and ML-class forecasts ($\tau>0.9$, VCR >> 99%) despite similar nominal performance in stable market conditions. Formally, RI = 0 across all five benchmarks, confirming that the misalignment is quantitative rather than ordinal.
The explanation follows directly from the structural argument of Section 3.2. Persistence imports yesterday’s intraday shape into today. When the realized intraday ordering changes, the forecast provides effectively no useful rank information for the current day’s DP problem, regardless of how close its absolute values are to realized prices. The ML forecast, by contrast, captures the shape of the intraday price curve with sufficient fidelity to preserve the ranking of profitable trading opportunities, even when individual price levels are imprecise.
Within the tau-sufficiency region, further forecast improvements yield negligible additional value. The Hybrid 8h forecast achieves a $\tau_{K}$ of 0.948 and a VCR of 99.7%—an improvement of only 0.5 percentage points over the ML 1-step forecast despite substantially lower MAE. For a battery generating approximately €1,750 per day, this corresponds to less than €9 per day of additional value from a more accurate forecast. The practical implication is direct: once a forecast achieves tau-sufficiency, investment in further accuracy improvements is economically unjustified for intraday dispatch.
The threshold is robust across market conditions. Repeating the synthetic experiment separately for normal years (2020–2021, 2024–2025) and crisis years (2022–2023), the tau-sufficiency region remains $[0.85,0.95]$ in both subsamples. Splitting the out-of-sample weeks by XBID volatility, the $\tau_{K}$ of ML-based forecasts remains stable at 0.929–0.934, comfortably above the threshold in both high- and low-volatility regimes. Persistence VCR, by contrast, drops by 13.9 percentage points in high-volatility weeks relative to low-volatility weeks, confirming that the MAE advantage of persistence is most misleading precisely when arbitrage opportunities are largest. Having established the forecast quality threshold for intraday dispatch, we turn to the capacity allocation decision that determines how much XBID budget is available in the first place.
4.2 FCR Dominance, Attribution, and Robustness Frontier
Table 3 reports revenue and VCR for the six configurations of the ablation study over the 2023–2025 out-of-sample period. The results reveal a clear hierarchy of value sources. The naive DA spread strategy captures only 30.9% of oracle value. Adding Layer 3 DP dispatch to a standalone battery with no reserve obligations achieves $\text{VCR}=100\%$ relative to its own XBID oracle, confirming that the Rolling Intrinsic algorithm is near-optimal given its SoC budget. The dominant value driver is Layer 1 reserve allocation: introducing FCR and aFRR participation raises average daily revenue from €492 to €1,696, reflecting the dominant contribution of reserve capacity payments in the German market during 2023–2025.
Table 3: Revenue attribution by system configuration (2023–2025, DE, out-of-sample). Each row adds one decision component. VCR is computed relative to a configuration-specific oracle (perfect foresight under identical market participation); no single oracle row is shown because each configuration uses a distinct upper bound. The oracle revenue for the full multi-market system (Layer 1 HMM+SAC) is €2,573/day. The HMM and SAC components operate as an integrated adaptive bidding system; no separate ablation between HMM and SAC is available.
[…]
Naive DA spread 492 46.8 0
DP only (Layer 3) 1,432 100.0 +53.2 398
Layer 2 ++ 3 †† 565 53.7 36
Layer 1 static (Q40) 1,696 64.3 64
Layer 1 HMM+SAC (full system) 1,753 68.1 ++3.8 95
Reserve only (static) 1,131 54.5 0
VCR relative to oracle_xbid_only (standalone XBID battery).
Revenue in EUR/day (mean 2023–2025). CVaR 5% = negative weekly revenue at 5th percentile.
††Shown for completeness; not included in the Figure 2 ablation chain.
The FCR allocation is endogenously fixed at 5 MW in all 157 out-of-sample weeks, reflecting the PICASSO buffer constraint, which limits maximum FCR participation to 50% of power capacity for a 1-hour battery under worst-case activation assumptions. The optimizer saturates FCR in every week, confirming that FCR capacity is consistently binding in the modeled system. The remaining 5 MW is allocated dynamically between aFRR (mean 0.65 MW, max 3.0 MW) and XBID (mean 4.38 MW). Removing the minimum XBID participation floor yields negligible revenue impact (+0.26% over three years), confirming that the result is not driven by the participation constraint. [^1] Without the floor, the optimizer allocates a minimum of 0.7 MW to XBID in all weeks, never choosing zero intraday participation. The floor is binding in 34 of 157 weeks (21.7%), with a revenue impact of €8k over three years.
[…]
The HMM+SAC adaptive bidding system—in which the HMM identifies the market regime and the SAC agent selects the optimal bid percentile for that regime—contributes €62,547 over three years relative to the static $Q_{40}$ strategy (+3.4%), driven by FCR acceptance rate improvement from 34% to 58% in 2023. The static $Q_{40}$ strategy, calibrated on the 2022 energy crisis, submits bids too high for the post-crisis normalization of 2023. The HMM identifies the regime shift and adjusts the bid percentile downward, recovering acceptance rate at a lower per-MW price. The revenue decomposition confirms the mechanism: FCR revenue increases by +14.8% under HMM adaptation, while aFRR ++ revenue declines by 31.5% as capacity is redirected toward FCR. The net effect is positive because the FCR gain exceeds the aFRR loss in absolute terms.
Figure 2: Value Capture Ratio (VCR) for each system configuration in the ablation study (out-of-sample 2023–2025, DE). Each configuration adds one decision layer. FCR/aFRR reserve layers contribute the majority of revenue; the XBID Layer 3 captures 15–18% of standalone oracle value due to SoC constraints from concurrent reserve obligations.
[…]
Table 4: Robustness frontier for the zero-missed-allocation result. The p90–p10 spread is used as the primary robust measure. The system operates at 0.86 cycles/day (measured), well below the frontier at 2.0 cycles/day.
Scenario Cycles/day Spread FCR/XBID Missed (%)
[…]
Measured from Layer 3 backtest (2023–2025). Trimodal distribution: modes at
0.44, 0.73, 1.03 cycles/day (see Section 4.2).
These findings jointly characterize the role of forecast quality in multi-market BESS systems. Because FCR capacity is saturated by the PICASSO buffer constraint and the residual intraday budget is small relative to total system capacity, the capacity allocation decision of Layer 1 is the primary driver of total revenue. The tau-sufficiency result of Section 4.1 completes the picture: for the intraday budget that is available, forecasts above $\tau^{*}\approx 0.85$ already capture near-oracle value, making further forecast improvements of limited incremental value for intraday dispatch decisions.
4.3 The Swiss Hydro Natural Experiment
To test the generality of our findings beyond the German market, we turn to the Swiss balancing market, which provides a natural experiment for identifying the role of exogenous structural factors in BESS revenue. Unlike the German market, where revenue is dominated by FCR capacity payments that are relatively stable across seasons, the Swiss SRL- market exhibits extreme price variability driven by hydrological conditions. When reservoir levels are anomalously high relative to seasonal norms, the TSO requires additional downward flexibility to absorb surplus hydraulic generation, driving SRL- prices sharply higher.
To identify this mechanism empirically, we construct a seasonal anomaly z-score from weekly Swiss reservoir level data, adjusting for the strong seasonal pattern in Swiss hydrology. The running variable is orthogonal to seasonality by construction, isolating the exogenous variation in reservoir conditions from regular winter drawdown and summer refill cycles.
Table 5 reports BESS revenue by hydrological regime. During high-hydro weeks ($z>0.7$), weekly SRL- revenue averages €20,022/MW, compared to €2,480/MW during low-hydro weeks (z<0.8z<-0.8) — a difference of $708\%$. The association is statistically significant: a one-standard-deviation increase in the anomaly is associated with approximately €6,400 of additional weekly revenue per MW (OLS, $p=0.0005$), and the SRL- price difference between HIGH and LOW regimes is itself highly significant (Spearman $p=0.0002$). The explanatory power is limited ($R^{2}=0.043$), as expected given the multiple drivers of weekly revenue.
Table 5: Swiss balancing market revenue by hydrological regime (2020–2025). The running variable is the hydrological anomaly z-score (seasonally adjusted). HIGH_HYDRO corresponds to $z>0.7$ standard deviations above the seasonal mean.
Regime N (weeks) SRL- price SRL- rev./week DA price
(EUR/MW/h) (EUR/MW) (EUR/MWh)
LOW_HYDRO (z<0.8z<-0.8) 78 16.2 2,480 108.2
MEDIUM 157 16.7 4,374 110.2
HIGH_HYDRO ($z>0.7$) 78 63.7 20,022 149.1
[…]
OLS total revenue: coeff =6,369=6{,}369 EUR/SD, $R^{2}=0.043$, $p=0.0005$.
The low $R^{2}$ is expected. Weekly BESS revenue depends on multiple drivers — DA prices, gas prices, aFRR dynamics — of which hydrological anomaly is only one. High statistical significance combined with low explanatory power is therefore consistent with a real but partial effect, as is standard in empirical market studies using macroeconomic running variables.
The lead-lag structure of the relationship provides additional evidence supporting the proposed mechanism. Spearman correlation between the hydro anomaly and SRL- prices is strongest at a lag of two to four weeks, suggesting that reservoir conditions are observable in advance and that price effects materialize with a delay as the TSO adjusts procurement volumes. This lead-lag structure is consistent with exogenous hydrological variation driving the price signal rather than the reverse.
Figure 3: Relationship between Swiss hydrological anomaly and SRL- balancing market revenue (2020–2025). Left: time series of hydrological anomaly z-score (blue) and SRL-dn capacity prices (green). Right: scatter plot of seasonal anomaly z-score vs. weekly BESS revenue from SRL-, with OLS fit ($R^{2}=0.043$, $p=0.0005$). The 2022–2023 storage scarcity crisis (Speicherkrise) is annotated.
We interpret these results as evidence consistent with an exogenous hydrological shock, without making a strict causal claim. The identification strategy relies on the orthogonality of the seasonal anomaly to regular market dynamics; potential confounders include correlated European weather events that simultaneously affect both reservoir levels and wholesale electricity prices. The finding nonetheless provides the first empirical documentation of the hydro-balancing market transmission mechanism in Switzerland, with direct implications for BESS operators in hydro-dependent markets such as Austria, Norway, and Iberia, where similar structural drivers may shape balancing market revenues.
5 Discussion
The three empirical findings of Section 4— the tau-sufficiency threshold, the FCR dominance under reserve constraints, and the Swiss hydro transmission mechanism—jointly reframe how BESS operators should evaluate forecast quality and allocate system capacity. We discuss their theoretical and practical implications in turn.
5.1 Theoretical Implication
The tau-sufficiency result has a theoretical implication that extends beyond battery storage applications. The DP Rolling Intrinsic algorithm is a specific instance of a broader class of optimization problems in which the objective function depends on the relative ordering of input parameters rather than their absolute values. A broad class of such problems—including certain forms of inventory management, scheduling, and trading—share this ordinal structure: the optimal policy is determined by the ranking of future states, not their precise magnitude. For any such problem, standard pointwise error metrics are misaligned with decision quality by construction, and rank correlation metrics are more appropriate evaluation criteria.
This connects to the emerging literature on decision-focused learning (Smets et al., 2024; Vanderschueren et al., 2022), which proposes retraining forecast models to minimize downstream decision cost rather than statistical error. While decision-focused learning has been shown to improve out-of-sample profits for energy storage systems, albeit at the cost of significant computational complexity (Smets et al., 2024), this approach integrates the downstream objective into model training rather than characterizing the decision structure itself. Our finding is complementary but distinct: we show that for problems with an ordinal decision structure, the alignment between statistical metrics and decision quality can be empirically characterized by a single threshold in rank correlation. This threshold is a property of the decision problem, not of the forecasting method, and is therefore stable across forecast architectures. For tau-sufficient forecasts, decision-focused retraining may offer limited additional value—a result that simplifies the operational pipeline for BESS operators. Consistent with Vanderschueren et al. (2022), who show empirically that decision-making strategies have a larger impact on cost performance than the objective function used for training, our results highlight that improvements in the decision framework can outweigh gains from modifying the predictive model itself.
5.2 Practical Implications for BESS Operators
The findings translate into three concrete operational recommendations.
First, forecast evaluation for intraday dispatch should prioritize rank-based metrics over pointwise error metrics such as MAE or RMSE. The relevant question for a BESS operator is not “what is my forecast error?” but “does my forecast preserve the ranking of intraday price opportunities?” In our setting, forecasts achieving $\tau_{K}>0.85$ are tau-sufficient: further investment in forecast accuracy will not materially improve dispatch revenue.
Second, investment in forecast improvement should prioritize Layer 1 over Layer 3. Our attribution results show that the capacity allocation decision dominates total revenue, while intraday dispatch quality is a second-order effect once tau-sufficiency is achieved. Improving the cross-market spread forecast that informs Layer 1 allocation—particularly during regime transitions when the static $Q_{40}$ strategy underperforms—has higher expected return than improving the intraday price forecast.
Third, operators in hydro-dependent markets should incorporate reservoir level monitoring into their bid strategy. The Swiss evidence shows that hydrological anomalies are observable two to four weeks in advance and are associated with SRL- price spikes that can increase weekly revenue by more than 700%. A simple regime-switching strategy that raises bid percentiles during high-hydro periods—analogous to the HMM mechanism applied to FCR in the German market—could capture a significant fraction of this opportunity.
5.3 Market Regime and Heterogeneity
The findings should be interpreted in the context of the German and Swiss markets over 2020–2025. Two sources of heterogeneity deserve attention.
The FCR dominance result depends on the current structure of the German balancing market, where BESS penetration in FCR has grown rapidly but has not yet saturated to the point of driving capacity prices to zero. As BESS capacity continues to expand, FCR prices may decline as BESS penetration increases, compressing the FCR/XBID revenue ratio. The tau-sufficiency result, by contrast, is structural: it depends on the ordinal nature of the DP dispatch problem, not on the relative profitability of FCR and XBID, and will remain valid as market conditions evolve.
The Swiss hydro finding highlights the importance of market-specific structural drivers that are absent from German-focused models. Similar mechanisms may operate in other hydro-dependent European markets—Austria, Norway, and Iberia—where hydroelectric generation plays a dominant role in reserve provision. Operators and researchers working in these markets should account for hydrological conditions as a primary revenue driver for downward balancing services, rather than treating balancing prices as purely driven by gas and demand dynamics.
5.4 Limitations and Scope
Three limitations of this study should be acknowledged.
First, aFRR energy activations are modeled via expected value and SoC buffer constraints rather than fully dynamic real-time simulation. This simplification is expected to be conservative for revenue estimation: the PICASSO buffer is calibrated to worst-case activation scenarios, so any dynamic activation exceeding the buffer would reduce available XBID capacity and bias revenue estimates downward rather than upward. However, the dynamic coupling between aFRR activations and XBID decisions is not fully modeled. In practice, concurrent activations reduce XBID capacity in real time, creating interactions that the layer-separated architecture does not capture. This may result in occasional infeasibility in high-activation periods that the static buffer approach does not detect.
[…]
Third, the tau-sufficiency threshold $\tau^{*}\approx 0.85$$0.95$ is estimated on German XBID data for 2023–2025. While the existence of a threshold is a structural property of ordinal decision problems, its precise numerical value may vary across markets with different price dynamics, liquidity structures, or trading horizons. Validation on other European markets would refine the quantitative estimate of $\tau^{*}$ and strengthen the generalizability of the finding.
6 Conclusion
This paper examines the relationship between forecast accuracy and decision quality in multi-market battery storage optimization, using six years of real market data from Germany and Switzerland.
Three findings emerge. First, we identify an empirical tau-sufficiency region ($\tau^{*}\approx 0.85$$0.95$) in Kendall rank correlation beyond which value capture saturates near 100% of oracle revenue, with limited sensitivity to further improvements in absolute forecast error. The DP dispatch problem is inherently ordinal—it depends on the relative ranking of prices, not their absolute levels—and any forecast that sufficiently preserves this ranking induces a near-optimal policy. Standard accuracy metrics such as MAE fail to capture this property, leading to systematic misalignment between statistical evaluation and economic performance.
Second, in multi-market systems under reserve constraints, the capacity allocation decision of Layer 1 is the primary driver of total revenue. FCR capacity is endogenously saturated by the PICASSO buffer constraint in all observed weeks of our sample, making the Layer 1 allocation—not the intraday forecast—the economically dominant decision. The tau-sufficiency result completes the picture: for the intraday budget that remains after reserve obligations, standard ML forecasts already achieve tau-sufficiency, making further forecast improvement of limited incremental value for intraday dispatch.
[…]
Together, these results reframe the forecast evaluation problem for BESS operators. The relevant question is not “what is my MAE?” but “does my forecast achieve tau-sufficiency, and is my capacity allocation strategy adapted to the current market regime?” Rank correlation metrics and regime-aware bidding, rather than pointwise accuracy and static strategies, are the primary tools that deliver economic value in heterogeneous European balancing markets.
Appendix A HMM+SAC Adaptive Bidding: Technical Details
We implement a Hidden Markov Model following the standard formulation of Rabiner (1989), using four observable features computed weekly: the rolling mean FCR clearing price (4-week window), its standard deviation, the week-over-week price change, and the aFRR acceptance rate. These features are standardized before training. The HMM is trained via Baum-Welch on the full 2020–2022 in-sample period and applied in a walk-forward fashion from 2023 onward, with the state sequence re-estimated monthly.
The four latent states correspond to identifiable market regimes: normal (FCR $\approx$ 16 EUR/MW/h), crisis (FCR $\geq$ 20 EUR/MW/h), post-crisis normalization (FCR declining rapidly), and low-volatility (FCR stable at low levels). Regime assignment is obtained from the most likely latent state sequence under the fitted model, using the Viterbi algorithm.
The Soft Actor-Critic (SAC) agent (Haarnoja et al., 2018) receives the current HMM state as input and outputs a bid percentile $p\in[Q_{20},Q_{60}]$, trained to maximize realized weekly FCR revenue. Training uses 2020–2022 data with a replay buffer of 52 weeks and a discount factor $\gamma=0.95$. The policy is fixed from 2023 onward (no online learning in the out-of-sample period). Validation on a held-out 2022 H2 window confirmed that the policy improves FCR acceptance rate relative to the static $Q_{40}$ strategy in all four regime states, with the largest gain in the post-crisis normalization state (acceptance rate: 34% $\to$ 58%). Out-of-sample performance is stable across all four regimes, with no evidence of overfitting in the held-out 2022 H2 validation window.
The HMM +SAC system adds €62,547 over three out-of-sample years relative to the static $Q_{40}$ baseline (+3.4%), with the improvement concentrated in the 2023 regime transition period. The module is modular and can be replaced with any regime-detection mechanism without affecting the Layer 1 LP structure or the Layer 3 DP dispatch.
Acknowledgements
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Declaration of Generative AI and AI-Assisted Technologies in the Manuscript Preparation Process
During the preparation of this work the author used Claude (Anthropic) to assist with manuscript drafting, literature positioning, and implementation. The author reviewed and edited all content, verified all numerical results against the underlying codebase and data, and takes full responsibility for the content of the published article.